A hiker hikes up a mountain 1mph slower than she hikes down the mountain. If it takes her 1.5 hours to hike up the mountain, and only 1 hour to hike down it, how fast does she move each direction?
Answer by checkley77(12569)

D=RT
D UP=(X-1)1.5
D DOWN=X*1
x=(x-1)1.5
X=1.5X-1.5
X-1.5X=-1.5
-.5X=-1.5
X=-1.5/-.5
X=3 MPH DOWN THE MY.
3-1=2 MPH UP THE MT.
PROOF:
3=(3-1)1.5
3=2*1.5
3=3